OpenStudy (anonymous):

5/7 of n is 7/3 .. what is n?

OpenStudy (acxbox22):

from that info we can set up an equation 5/7n=7/3 now can u solve for n?

OpenStudy (anonymous):

I got 3.26666666667. But i need it as a fraction

OpenStudy (anonymous):

So 15/21n = 49/21

OpenStudy (acxbox22):

no you dont need common denominators 5/7n=7/3 multiply both sides by the reciprocal of 5/7 to find n

OpenStudy (anonymous):

49/15 or 3 4/15?